OpenStudy (anonymous):

John is twice as old as Mary. The sum of their ages is 21. How old is Mary? Let J = John's age and M = Mary's age. Select the system equations that represents the problem. J - 2M = 0 M - 2J = 0 J + M = 21 J = M + 2 b,c ?

OpenStudy (anonymous):

....its multiple choice

OpenStudy (zpupster):

c is right

OpenStudy (zpupster):

what else??

OpenStudy (anonymous):

i dont know i was just wondering which ones are correct......so just c

OpenStudy (wolf1728):

Also a

OpenStudy (anonymous):

no

OpenStudy (zpupster):

no a is correct also J = 2M then subtract 2M from both sides

OpenStudy (wolf1728):

From the problem we know: X) J = 2*M Y) J + M = 21 Substituting 2*M into Y Y) 2*M + M =21 M = 7 J = 14 So these equations are correct a) J - 2M = 0 c) J + M = 21

OpenStudy (anonymous):

thanks guuys

OpenStudy (zpupster):

good job!!!

OpenStudy (wolf1728):

u r welcome